OpenStudy (anonymous):

help plz Write an equation of a line perpendicular to y = 1/2x - 5 in slope-intercept form that passes through the point (2, -7). y=1/2 x - 7 y= 2x + 2 y= -2x - 5 y= -2x - 3

OpenStudy (zale101):

So, you are given an equation to determine a perpendicular line with a given point. From the given equation y = 1/2x - 5 you identify the slope. Convert the original slope into a perpendicular line slope, and by that, you take the negative reciprocal of it. \(\Large m=-\frac{1}{m}\) Place the new perpendicular slope and the given point (2,-7)=(x1,y1) into the point slope formula \(y-y_1=-\frac{1}{m}(x-x_1)\)

OpenStudy (anonymous):

so y- (-7) = -1/m(x-2) where do i go from here @Zale101

OpenStudy (anonymous):

@Hero could you please help me finish this